i need help solving the equations and checking your solution
4/3x-6/x=-4

Answer by checkley77(12844)

4/3x-6/x=-4
(4*x-6*3x)/3x^2=-4
(4x-18x)/3x^2=-4
-14x/3x^2=-4
-4*3x^2=-14x
-12x^2+14x=0
-2x(6x-7)=0
6x-7=0
x=7/6 answer.